Aiello, Inc. had the following inventory in fiscal 2016. The company uses the FIFO method of accounting for inventory. Beginning Inventory, January 1, 2016: 130 units @ $15.00 Purchase 200 units @ $18.00 Purchase 50 units @ $13.50 Purchase 110 units @ $15.75 Ending Inventory, December 31, 2016: 120 units The company’s cost of goods sold for fiscal 2016 is: Select one: A. $6,090.00 B. $6,045.00 C. $6,157.50 D. $5,305.75 E. None of the above
Answer #1

Units sold = (130+200+50+110) - 120 = 370 units

Cost of Goods Sold under Fifo

= (130*15) + (200*18) + (40*13.50)

= 6090

Option A is the answer